Solve for n.
R=Pi/1-(1+i)^-n

Answer by stanbon(75887)

Solve for n.
R=Pi/1-(1+i)^-n
----
1 - (1+i)^-n = pi/R
-----
(1+i)^-n = 1-[pi/R]
-----
(1+i)^n = 1/[1-(pi/R)]
n*log(1+i) = -log[1-(pi/R)
----
n = -log[(R-pi)/R]/log(1+i)
=================================
Cheers,
Stan H.
===================